The perimeter if a rectangular lawn is 56.If the area of the same lawn is 187, find the length and breadth of the same lawn.
Answer by Alan3354(69443)

The perimeter if a rectangular lawn is 56.If the area of the same lawn is 187, find the length and breadth of the same lawn.
============
P = 2L + 2W = 56
L + W = 28
L*W = 187
--------
Find a pair of factors of 187 with a sum of 28.